# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ 3.2.2 GRR-MLMC
The Russian roulette (RR) estimator [44] is also employed to estimate the sum of an infinite series, wherein the evaluation of any term in the series only demands a finite amount of computation. This estimator relies on randomized truncation and assigns a higher weight to each term to accommodate the possibility of not computing them. The query of the gradient of RR-MLMC estimator is given by
@@
Motivated by [43], in order to trade lower variance with higher cost based on the conventional RR-MLMC, one way is to ensure that the first ${\underline{m}}$ terms of the infinite series are always computed, where this ${\underline{m}}$ is called the base level. In this case, the random index $L$ is set to have a lower bound ${\underline{m}}$, i.e., $\mathbb{P}(L\geq{\underline{m}})=0$, implying $p_{j}=1$ for all $j\leq{\underline{m}}$. We call this modified estimator as generalized Russian roulette (GRR) estimator. When ${\underline{m}}=0$, this degenerates to the conventional RR-MLMC. The associated quaery for GRR-MLMC estimator of the gradient is
@@
In this case, the probability mass function of $L$ is chosen as $\mathbb{P}(L={\underline{m}})=1-\sum_{\ell>{\underline{m}}}w_{\ell}$ and $\mathbb{P}(L=\ell)=w_{\ell}$ for $\ell>{\underline{m}}$, where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}})$. This implies that $V_{\mathrm{GRR}}$ is also unbiased since all $p_{j}>0$,
Similarly, we define the query for GRR estimator of the loss function:
@@
$U_{\mathrm{GRR}}$ and the following theorem is established aiming at evaluating their order.
Theorem 3.6.
Under the setting of Theorems 3.2 and 3.3, let $L\sim\mathrm{Ge}(p,{\underline{m}})$ with $p=1-2^{-\alpha}$ and $1<\alpha<\min\left(r_{1},r_{2}\right)$, where $r_{1}$ and $r_{2}$ are from Theorems 3.2 and 3.3, and the base level ${\underline{m}}\geq 1$, we have
$\mathrm{Var}\left[U_{\mathrm{GRR}}\right]$ $\leq A\left(\frac{1}{2^{{\underline{m}}}}+\frac{2^{(\alpha-r_{1})({\underline{m}}+1)}}{1-2^{\alpha-r_{1}}}\right),$
$\mathrm{Var}\left[V_{\mathrm{GRR}}\right]$ $\leq B\left(\frac{1}{2^{{\underline{m}}}}+\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right),$
$\mathrm{Cost}_{\mathrm{GRR}}$ $\propto M_{0}2^{{\underline{m}}}+M_{0}\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}},$
where the constants $A,B$ are independent of $\alpha,\ r_{1},\ r_{2},\ \underline{m}$.
The proof of this theorem is detailed in Appendix A.3.
@@ 3.3 Truncated MLMC for APT
When treating $\nabla_{\phi}\mathcal{L}(\phi)$ as the summation of the infinite random variable sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, we notice that $\Delta\rho_{\phi,0}=\rho_{\phi,0}$ is just the nested APT estimator with a bias of order $O(1/M_{0})$. For $\Delta\rho_{\phi,\ell}$ with $\ell\geq 1$, it actually contributes to the reduction of bias while simultaneously increasing the variance and this additional variance grows with $\ell$. An empirical demonstration of the additional variance for $\Delta\rho_{\phi,\ell}$ is shown in Figure 2.
@@
We observe that the truncated RU-MLMC method enjoys a more stable loss compared to the original one, indicating that truncation helps reducing variance. Therefore, in order to mitigate the variance of the RU-MLMC, we propose to truncate the distribution of $L$ by setting the largest value of $L$ as $\overline{m}$. The truncation leads to a bias of order $O(1/M_{{\overline{m}}})$. Following [33], we extend the truncated idea to other unbiased MLMC methods.
@@ 3.3.1 TGRR-MLMC
TGRR-MLMC stands for the truncated version of GRR-MLMC. Let ${\overline{m}}\geq{\underline{m}}$ be a truncated level, and let $L$ be a nonnegative integer-value random index taking values in $\{{\underline{m}},\dots,{\overline{m}}\}$. The query for the estimation of the gradient of TGRR-MLMC is then defined as
$V_{\mathrm{TGRR}}=\rho_{\phi,M_{{\underline{m}}}}+\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\rho_{\phi,j}}{p_{j}},$ (24)
@@
We should note that if ${\overline{m}}={\underline{m}}=m$, TGRR-MLMC degenerates to nested APT with the inner sample size $M_{m}=M_{0}2^{m}$. For any truncated method with truncated level $\overline{m}$, they all share the bias of order $\mathcal{O}(M_{\overline{m}})$ based on (45).
@@
Input: Prior $p(\theta)$, implicit simulator model $p(x|\theta)$, neural network $q_{F(x,\phi)}(\theta)$, $1<\alpha<\min(r_{1},r_{2})$, $M_{0},M_{\ell}=M_{0}2^{\ell}$, truncated level ${\overline{m}}$, base level ${\underline{m}}$, total round $R$, optimizer $g(\cdot,\cdot)$, batch size $B$, outer sample size $N$, $t=0$, dataset $\mathcal{D}_{\mathrm{out}}$ for outer samples, dataset $\mathcal{D}_{\mathrm{in}}(x)$ for the inner samples of $x$, $\tilde{p}_{1}(\theta)=p(\theta)$
1 for $k$ in $\{1,2,\cdots,R\}$ do Generate $\{(\theta_{i},x_{i})\}_{i=1}^{N}$ from $\tilde{p}_{k}(\theta)p(x|\theta)$ and level $\{\ell_{i}\}_{i=1}^{N}$ from $\mathrm{Ge}(p,{\underline{m}},{\overline{m}})$ Update dataset $\mathcal{D}_{\mathrm{out}}\leftarrow\{(\theta_{i},x_{i},\ell_{i})\}_{i=1}^{N}\cup\mathcal{D}_{\mathrm{out}}$
2 if $k=1$  then repeat Generate $\{(\theta_{i},x_{i})\}_{i=1}^{B}\sim\mathcal{D}_{\mathrm{out}}$ Update parameter $\phi_{t+1}\leftarrow g\left(\phi_{t},-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\phi}\log q_{F(x_{i},\phi_{t})}(\theta_{i})\right)$ $t\leftarrow t+1$ until $\phi_{t}$ converged
3end if
4if $k\geq 2$ then repeat Generate $\{(\theta_{i^{\prime}},x_{i^{\prime}})\}_{i=1}^{B}\sim\mathcal{D}_{\mathrm{out}}$
5 for $i^{\prime}$ in $\{1,2,\cdots,B\}$ do Generate $\{\theta^{\prime}_{j}\}_{j=1}^{M_{{\underline{m}}}-1}$ from $\tilde{p}_{k}(\theta^{\prime})$ $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=1}^{M_{{\underline{m}}}-1}\cup\{\theta_{i^{\prime}}\}$ and compute $\rho^{(i^{\prime})}_{\phi_{t},M_{m}}$ with $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})$ for $\ell^{\prime}$ in $\{\{{\underline{m}}+1,{\underline{m}}+2,\cdots,\ell_{i}\}$} do $\mathcal{D}^{(a)}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=1}^{M_{\ell^{\prime}-1}-1}\cup\{\theta_{i^{\prime}}\}$, $\mathcal{D}^{(b)}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=M_{\ell^{\prime}-1}}^{M_{\ell^{\prime}}-1}\cup\{\theta_{i^{\prime}}\}$ $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\mathcal{D}^{(a)}_{\mathrm{in}}(x_{i^{\prime}})\cup\mathcal{D}^{(b)}_{\mathrm{in}}(x_{i^{\prime}})$ Compute $\Delta\rho^{(i^{\prime})}_{\phi_{t},\ell^{\prime}}$ with $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})$ using (19) end for
6Compute $V_{\mathrm{TGRR}}^{(i^{\prime})}$ using (24) end for
7Update parameter $\phi_{t+1}\leftarrow g\left(\phi_{t},-\frac{1}{B}\sum_{i^{\prime}=1}^{B}V_{\mathrm{TGRR}}^{(i^{\prime})}\right)$ $t\leftarrow t+1$ until $\phi_{t}$ converged
8end if
9Update proposal $\tilde{p}_{k+1}(\theta)\leftarrow q_{F(x_{o},\phi_{t})}(\theta)$ end for
Algorithm 1 Truncated Generalized Russian Roulett (TGRR)
@@ 4 Convergence results of SGD
In this section, we analyze the convergence of the methods proposed in this paper with some existing techniques. For simplicity, we only study the widely used SGD with constant step size $\gamma$ in this paper. Our result should extend to boarder cases. The parameter update procedure of SGD is presented as
$\phi_{t+1}=\phi_{t}-\gamma_{t}\nabla_{\phi}\hat{\mathcal{L}}(\phi_{t}),$
@@
where $b(\phi)$ and $\eta(\phi)$ denote the bias and the noise of gradient estimator $\nabla_{\phi}\hat{\mathcal{L}}(\phi)$ respectively.
@@
There exist constants $U_{b}<\infty$ and $U_{\eta}<\infty$ such that
$\underset{\phi\in\Phi}{\sup}\left\|b(\phi)\right\|_{2}^{2}$ $=\underset{\phi\in\Phi}{\sup}\left\|\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]-\nabla_{\phi}\mathcal{L}(\phi)\right\|^{2}_{2}\leq U_{b},$
$\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left\|\eta(\phi)\right\|_{2}^{2}\right]$ $=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left\|\nabla_{\phi}\hat{\mathcal{L}}(\phi)-\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]\right\|_{2}^{2}\right]\leq U_{\eta}.$
Following the similar procedure to derive (45), we find that the order of $U_{b}$ for nested APT is $\mathcal{O}(1/M)$, while $\mathcal{O}(1/(M_{0})^{\overline{m}})$ for TGRR. Intuitively speaking, when it is set that $M_{0}^{\overline{m}}=M$ and $\underline{m}<\overline{m}$, $M$ inner samples are always required for each outer sample in nested APT, while the inner sample size for TGRR-MLMC is determined by the distribution $\mathrm{Ge}(p,\underline{m},\overline{m})$ whose maximal output is $M_{0}^{\overline{m}}=M$. Therefore, this truncated MLMC method takes less cost than nested APT to achieve the same level of bias.
We then focus on studying how $U_{b}$ and $U_{\eta}$ would affect the optimal gap at any $t>0$, which is defined as
@@
where $\phi^{*}=\underset{\phi\in\Phi}{\arg\min}\,\mathcal{L}(\phi)$. Moving forward, we introduce the following two commonly used assumptions when analyzing the convergence of SGD, as outlined in [6].
Assumption 4.2.
The objective function $\mathcal{L}(\phi)$ is differentiable and there exists a constant $K$ such that for every $\phi,\phi^{\prime}$
$\mathcal{L}(\phi)\leq\mathcal{L}(\phi^{\prime})+\nabla_{\phi}\mathcal{L}(\phi)^{T}(\phi-\phi^{\prime})+\frac{K}{2}\left\|\phi-\phi^{\prime}\right\|^{2}_{2}.$
Assumption 4.3.
The objective function $\mathcal{L}(\phi)$ is differentiable and there exists $\mu>0$ such that
$\left\|\nabla\mathcal{L}(\phi)\right\|_{2}^{2}\geq 2\mu(\mathcal{L}(\phi)-\mathcal{L}(\phi^{*})).$
In situations where a fixed number of iterations are exclusively employed for parameter updates, the evaluation of the impact of variance and bias on the upper bound of the optimal gap is undertaken through the application of the theorem presented below, following [1].
@@
Under Assumptions 4.2 and 4.3 and the constant step size $\gamma\leq\min\left\{1/K,1/\mu\right\}$, for a fixed $T$ steps of SGD, the upper bound of the optimal gap (27) is given as
@@
Since $G_{0}$ depends solely on the initialization of $\phi$, our focus lies on the last two terms of (28), where $U_{b}$ and $U_{\eta}$ have the same impact on the optimal gap $G_{T}$. This implies that unbiased methods with large variance may ultimately achieve the same result as those biased ones at the same iteration step $T$. Hence in cases where variance dominates bias, methods with smaller variance are preferable.
On the other hand, when a sufficiently large iteration step $T$ is given and proper learning rate $\gamma$ is chosen, the upper bound of the optimal gap can be smaller than any given $\epsilon>0$ except for the parts of bias, which is examined in the following corollary.
Corollary 4.5.
Suppose that Assumptions 4.2 and 4.3 are satisfied and $\gamma=\min\left\{1/K,1/\mu,\epsilon\mu/(K\eta)\right\},\,T=\max\left\{(K/\mu)\ln(2G_{0}/\epsilon),K\eta/(\epsilon\mu^{2})\ln(2G_{0}{\epsilon})\right\}$. Then the optimal gap satisfies $G_{T}\leq\epsilon+U_{b}/(2\mu)$ for any $\epsilon>0$.
@@
It follows directly from (28) by scaling $\gamma$ and $T$. ∎
In this case, unbiased methods are always favorable, as they are able to converge to an arbitrarily small neighborhood of 0, while biased methods can only converge to a neighborhood of the $U_{b}$. Moreover, for biased method with $U_{b}\gg U_{\eta}$, the optimal gap for $t_{1}\gg t_{2}$ tends to be $G_{t_{1}}\approx G_{t_{2}}$, indicating that increasing the iteration step can be futile in attempting to reduce the optimal gap.

### v2

@@ 3.2.2 GRR-MLMC
The Russian Roulette (RR) estimator [43] is also employed to estimate the sum of an infinite series, wherein the evaluation of any term in the series only demands a finite amount of computation. This estimator relies on randomized truncation and assigns a higher weight to each term to accommodate the possibility of not computing them. The query of the gradient of the RR-MLMC estimator is given by
@@
Motivated by [42], to trade lower variance with higher cost based on the conventional RR-MLMC, one way is to ensure that the first ${\underline{m}}$ terms of the infinite series are always computed, where this ${\underline{m}}$ is called the base level. In this case, the random index $L$ is set to have a lower bound ${\underline{m}}$, i.e., $\mathbb{P}(L\geq{\underline{m}})=1$, implying $p_{j}=1$ for all $j\leq{\underline{m}}$. We call this modified estimator the generalized Russian roulette (GRR) estimator. When ${\underline{m}}=0$, this degenerates to the conventional RR-MLMC. The associated query for the GRR-MLMC estimator of the gradient is
@@
In this case, the probability mass function of $L$ is chosen as $\mathbb{P}(L={\underline{m}})=1-\sum_{\ell>{\underline{m}}}w_{\ell}$ and $\mathbb{P}(L=\ell)=w_{\ell}$ for $\ell>{\underline{m}}$, where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}})$.
Similarly, we define the query for the GRR estimator of the loss function:
@@
We next provide an upper bound for the variance of the loss $U_{\mathrm{GRR}}$. It is sufficient to bound $\mathrm{Var}\left[\psi_{\phi,M_{{\underline{m}}}}\right]$ and $\mathrm{Var}\left[\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\psi_{\phi,j}}{p_{j}}\right]$. Directly applying Theorem 3.1 and the law of total expectation, we have
$\mathrm{Var}\left[\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\psi_{\phi,j}}{p_{j}}\right]$ $\leq\mathbb{E}\left[\left(\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\psi_{\phi,j}}{p_{j}}\right)^{2}\right]\leq\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\mathbb{E}\left[\left(\sum_{j={\underline{m}}+1}^{\ell}\frac{\Delta\psi_{\phi,j}}{p_{j}}\right)^{2}\right]$
$\leq\sum_{\ell={\underline{m}}+1}^{\infty}(\ell-\underline{m})w_{\ell}\sum_{j={\underline{m}}+1}^{\ell}p_{j}^{-2}\mathbb{E}\left[\Delta\psi_{\phi,j}^{2}\right]$
$\leq B_{1}\sum_{\ell={\underline{m}}+1}^{\infty}(\ell-\underline{m})w_{\ell}\sum_{j={\underline{m}}+1}^{\ell}2^{(2\alpha-r_{1})j}$
$\leq\frac{B_{1}p2^{(2\alpha-r_{1})(\underline{m}+1)}}{1-2^{2\alpha-r_{1}}}\sum_{\ell={\underline{m}}+1}^{\infty}(\ell-\underline{m})2^{-\alpha\ell}(1-2^{(2\alpha-r_{1})(\ell-\underline{m})})$
$\leq\frac{B_{1}p2^{(\alpha-r_{1})\underline{m}}}{1-2^{-2\alpha+r_{1}}}\left(\sum_{\ell=1}^{\infty}\ell 2^{-(r_{1}-\alpha)\ell}-\sum_{\ell=1}^{\infty}\ell 2^{-\alpha\ell}\right)$
$\leq\frac{B_{1}p2^{(\alpha-r_{1})\underline{m}}}{1-2^{-2\alpha+r_{1}}}\left(\frac{2^{-(r_{1}-\alpha)}}{(1-2^{-(r_{1}-\alpha)})^{2}}-\frac{2^{-\alpha}}{(1-2^{-\alpha})^{2}}\right)$
$\leq\frac{B_{1}(1-2^{-\alpha})2^{(\alpha-r_{1})(\underline{m}+1)}}{(1-2^{-2\alpha+r_{1}})(1-2^{\alpha-r_{1}})^{2}},$
where $B_{1}$ represents the implied constant in the upper bound of the second moment in Theorem 3.1. Note that
$\mathrm{Var}\left[\psi_{\phi,M_{\underline{m}}}\right]\leq 2\mathrm{Var}\left[\log g_{\phi}(x,\theta)\right]+2\mathrm{Var}\left[\log\hat{Z}_{M_{\underline{m}}}(x,\phi)\right],$
and $\mathrm{Var}\left[\log\hat{Z}_{M_{\underline{m}}}(x,\phi)\right]\leq A_{1}2^{-\underline{m}}$ for a constant $A_{1}>0$, as shown in [53]. Finally,
$\mathrm{Var}\left[U_{\mathrm{GRR}}\right]\leq 2\mathrm{Var}\left[\log g_{\phi}(x,\theta)\right]+A_{1}2^{-\underline{m}}+\frac{2B_{1}(1-2^{-\alpha})2^{(\alpha-r_{1})(\underline{m}+1)}}{(1-2^{-2\alpha+r_{1}})(1-2^{\alpha-r_{1}})^{2}}.$
We can similarly bound $\mathrm{Var}\left[V_{\mathrm{GRR}}\right]$ by replacing $r_{1}$ with $r_{2}$ from Theorem 3.2, yielding
$\mathrm{Var}\left[V_{\mathrm{GRR}}\right]\leq 2\mathrm{Var}\left[\nabla_{\phi}\log g_{\phi}(x,\theta)\right]+A_{2}2^{-\underline{m}}+\frac{2B_{2}(1-2^{-\alpha})2^{(\alpha-r_{2})(\underline{m}+1)}}{(1-2^{-2\alpha+r_{2}})(1-2^{\alpha-r_{2}})^{2}},$
where $A_{2}>0,B_{2}>0$ are constants independently of $\underline{m},\alpha,r_{2}$.
For the cost, we have the following
$=M_{0}2^{\underline{m}}+M_{0}\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}}.$
As we can predict, a larger $\underline{m}$ leads to lower variances, but it increases the costs. On the other hand, for a fixed $\underline{m}$, a larger $\alpha$ leads to a lower cost, but it raises the variances.
@@ 3.3 Truncated MLMC for APT
When treating $\nabla_{\phi}\mathcal{L}(\phi)$ as the summation of the infinite random variable sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, we notice that $\Delta\rho_{\phi,0}=\rho_{\phi,0}$ is just the nested APT estimator with a bias of order $\mathcal{O}(1/M_{0})$. For $\Delta\rho_{\phi,\ell}$ with $\ell\geq 1$, it actually contributes to the reduction of bias while simultaneously increasing the variance and this additional variance grows with $\ell$. An empirical demonstration of the additional variance for $\Delta\rho_{\phi,\ell}$ is shown in Figure 1.
@@
We observe that the truncated RU-MLMC method enjoys a more stable loss compared to the original one, indicating that truncation helps reducing variance. Therefore, in order to mitigate the variance of the RU-MLMC, we propose to truncate the distribution of $L$ by setting the largest value of $L$ as $\overline{m}$. The truncation leads to a bias of order $\mathcal{O}(1/M_{{\overline{m}}})$. Following [32], we extend the truncated idea to other unbiased MLMC methods.
@@ 3.3.1 TGRR-MLMC
TGRR-MLMC stands for the truncated version of GR
R-MLMC. Let ${\overline{m}}\geq{\underline{m}}$ be a truncated level, and let $L$ be a nonnegative integer-value random index taking values in $\{{\underline{m}},\dots,{\overline{m}}\}$. The query for the estimation of the gradient of TGRR-MLMC is then defined as
$V_{\mathrm{TGRR}}:=\rho_{\phi,M_{{\underline{m}}}}+\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\rho_{\phi,j}}{p_{j}},$ (26)
@@
We should note that if ${\overline{m}}={\underline{m}}=m$, TGRR-MLMC degenerates to nested APT with the inner sample size $M_{m}=M_{0}2^{m}$. For any truncated method with truncated level $\overline{m}$, they all share the bias of order $\mathcal{O}(2^{-\overline{m}})$.
@@
Input: Prior $p(\theta)$, implicit simulator model $p(x|\theta)$, conditional density estimator $q_{F(x,\phi)}(\theta)$, $\alpha\in(1,\min(r_{1},r_{2}))$, $M_{0}$, $M_{\ell}=M_{0}2^{\ell}$, truncated level ${\overline{m}}$, base level ${\underline{m}}$, total round $R$, optimizer $\mathrm{opt}(\cdot,\cdot)$, batch size $B$, outer sample size $N$
1 Initialization: Set iteration step $t\leftarrow 0$ and initial network parameter $\phi_{0}$, initialize outer sample dataset $\mathcal{D}_{\mathrm{out}}\leftarrow\emptyset$, set proposal $\tilde{p}_{1}(\theta)\leftarrow p(\theta)$
2for $k$ in $\{1,2,\cdots,R\}$ do
3       Generate $\{(\theta_{i},x_{i})\}_{i=1}^{N}$ from $\tilde{p}_{k}(\theta)p(x|\theta)$
4      Update dataset $\mathcal{D}_{\mathrm{out}}\leftarrow\{(\theta_{i},x_{i})\}_{i=1}^{N}\cup\mathcal{D}_{\mathrm{out}}$
5      if $k=1$ then
6             repeat
7                   Sample $\{(\tilde{\theta}_{i},\tilde{x}_{i})\}_{i=1}^{B}$ from $\mathcal{D}_{\mathrm{out}}$ without replacement
8                  Update parameter $\phi_{t+1}\leftarrow\mathrm{opt}\left(\phi_{t},-\frac{1}{B}\sum_{i=1}^{B}\nabla_{\phi_{t}}\log q_{F(\tilde{x}_{i},\phi_{t})}(\tilde{\theta}_{i})\right)$
9                  $t\leftarrow t+1$
10            until $\phi_{t}$ converged;
11       end if
12
13      if $k\geq 2$ then
14             repeat
15                   Sample $\{(\tilde{\theta}_{i},\tilde{x}_{i})\}_{i=1}^{B}$ from $\mathcal{D}_{\mathrm{out}}$ without replacement
16                  Sample level $\{\ell_{i}\}_{i=1}^{B}$ from $\mathrm{Ge}(1-2^{-\alpha},{\underline{m}},{\overline{m}})$
17                  for $i$ in $\{1,2,\cdots,B\}$ do
18                         Sample $\{\theta^{\prime}_{j}\}_{j=1}^{M_{{\underline{m}}}}$ from those $\theta$ in $\mathcal{D}_{\mathrm{out}}$ without replacement
19                        Compute $\rho^{(i)}_{\phi_{t},M_{\underline{m}}}=-\nabla_{\phi_{t}}\log g_{\phi_{t}}(\tilde{x}_{i},\tilde{\theta}_{i})+\nabla_{\phi_{t}}\log\frac{1}{M_{\underline{m}}}\sum_{j=1}^{M_{\underline{m}}}g_{\phi_{t}}(\tilde{x}_{i},\theta^{\prime}_{j})$
20                        for $\ell$ in $\{{\underline{m}}+1,{\underline{m}}+2,\cdots,\ell_{i}$} do
21                               Sample $\{\theta^{\prime}_{j}\}_{j=1}^{M_{\ell}}$ from those $\theta$ in $\mathcal{D}_{\mathrm{out}}$ without replacement
22                              Compute $\Delta\rho^{(i)}_{\phi_{t},\ell}$ by (20) with $x=\tilde{x}_{i}$, $\phi=\phi_{t}$ and $\{\theta^{\prime}_{j}\}_{j=1}^{M_{\ell}}$
23                         end for
24                        Compute the gradient $V_{\mathrm{TGRR}}^{(i)}$ using (26)
25                   end for
26
27                  Update parameter $\phi_{t+1}\leftarrow\mathrm{opt}\left(\phi_{t},\frac{1}{B}\sum_{i=1}^{B}V_{\mathrm{TGRR}}^{(i)}\right)$
28                  $t\leftarrow t+1$
29            until $\phi_{t}$ converged;
30       end if
31      Update proposal $\tilde{p}_{k+1}(\theta)\leftarrow q_{F(x_{o},\phi_{t})}(\theta)$
32 end for
Algorithm 1 Truncated Generalized Russian Roulette (TGRR)
@@ 4 Convergence results of SGD
In this section, we analyze the convergence of the methods proposed in this paper with some existing techniques. For simplicity, we only study the widely used SGD with constant step size $\gamma$ in this paper. Our result should extend to border cases. The parameter update procedure of SGD is presented as
$\phi_{t+1}=\phi_{t}-\gamma\nabla_{\phi}\hat{\mathcal{L}}(\phi_{t}),$ (28)
@@
where $b(\phi)=\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]-\nabla_{\phi}\mathcal{L}(\phi)$ and $\eta(\phi)=\nabla_{\phi}\hat{\mathcal{L}}(\phi)-\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]$ denote the bias and the noise of gradient estimator $\nabla_{\phi}\hat{\mathcal{L}}(\phi)$, respectively.
@@
Assume that
$U_{b}$ $:=\underset{\phi\in\Phi}{\sup}\left\|b(\phi)\right\|_{2}^{2}<\infty,\quad U_{\eta}:=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left\|\eta(\phi)\right\|_{2}^{2}\right]<\infty.$
We focus on studying how $U_{b}$ and $U_{\eta}$ would affect the optimal gap at any $t>0$, which is defined as
@@
where $\phi^{*}=\underset{\phi\in\Phi}{\arg\min}\,\mathcal{L}(\phi)$. In situations where a fixed number of iterations is exclusively employed for parameter updates, the evaluation of the impact of variance and bias on the upper bound of the optimal gap is undertaken through the application of the theorem presented below, following [1].
@@
Assume that 4.1 is satisfied and there exist positive constants $K,\mu$ such that for every $\phi,\phi^{\prime}\in\Phi$,
$\mathcal{L}(\phi)$ $\leq\mathcal{L}(\phi^{\prime})+\nabla_{\phi}\mathcal{L}(\phi)^{T}(\phi-\phi^{\prime})+\frac{K}{2}\left\|\phi-\phi^{\prime}\right\|^{2}_{2},$ (31)
$\left\|\nabla_{\phi}\mathcal{L}(\phi)\right\|_{2}^{2}$ $\geq 2\mu(\mathcal{L}(\phi)-\mathcal{L}(\phi^{*})).$ (32)
If the SGD algorithm (28) has a step size $\gamma\leq\min(1/K,1/\mu)$, for a fixed $T$ iterations, the upper bound of the optimal gap (30) is given as
@@
If $\gamma=\min\left\{1/K,\epsilon\mu/(KU_{\eta})\right\}<1/\mu$ and
$T=\max\left\{(K/\mu)\ln(2G_{0}/\epsilon),KU_{\eta}/(\epsilon\mu^{2})\ln(2G_{0}/{\epsilon})\right\},$
then the optimal gap satisfies $G_{T}\leq\epsilon+U_{b}/(2\mu)$ for any $\epsilon>0$.
@@
This is a special case of Theorem 6 in [1] by taking $m=M=0$ there.
Since $G_{0}$ depends solely on the initialization of $\phi$, our focus lies on the last two terms of (33), where $U_{b}$ and $U_{\eta}$ have the same impact on the optimal gap $G_{T}$. This implies that unbiased methods with large variance may ultimately achieve the same result as those biased ones at the same iteration step $T$. Hence in cases where variance dominates bias, methods with smaller variance are preferable. On the other hand, when a sufficiently large iteration step $T$ is given and proper learning rate $\gamma$ is chosen, the upper bound of the optimal gap can be smaller than any given $\epsilon>0$ except for the bias, which is examined in the second part of theorem 4.2. In this case, unbiased methods are always favorable, as they can converge to an arbitrarily small neighborhood of 0, while biased methods can only converge to a neighborhood of the $U_{b}/(2\mu)$. Moreover, for biased methods with $U_{b}\gg U_{\eta}$, the optimal gap for $t_{1}\gg t_{2}$ tends to be $G_{t_{1}}\approx G_{t_{2}}$, indicating that increasing the iteration step can be futile in attempting to reduce the optimal gap.